Gabriela had a snack of 7 carrots and 9 Hershey kisses for 355 calories. Ciarra had 5 carrots and 12 Hershey kisses for 265 calories. How many calories are in one carrot? How many calories are in one Hershey kiss?
Answer by ankor@dixie-net.com(22740)

Gabriela had a snack of 7 carrots and 9 Hershey kisses for 355 calories.
Ciarra had 5 carrots and 12 Hershey kisses for 265 calories.
How many calories are in one carrot? How many calories are in one Hershey kiss?
;
let h = no. of calories in a Hershey kiss
let c = no. of calories in a carrot
:
write an equation for each statement
"Gabriela had a snack of 7 carrots and 9 Hershey kisses for 355 calories.
7c + 9h = 355
"Ciarra had 5 carrots and 12 Hershey kisses for 265 calories."
5c + 12h = 265
we will use elimination here, multiply the 1st equation by 5, the 2nd by 7
35c + 45h = 1775
35c + 84h = 1855
for subtraction it is better to rearrange it:
35c + 84h = 1855
35c + 45h = 1775
------------------subtraction eliminates c, find h
0 + 39h = 80
h = 80/39
h = 2.0513 ~ 2 calories per kiss
:
Use the first equation to find the no. of calories in the carrot
7c + 9(2) = 355
7c = 355 - 18
7c = 337
c = 337/7
c = 48.143 ~ 48 calories per carrot
:
:
Check these solutions in the 2nd equation
5(48) + 12(2) = 265
240 + 24 = 264, close enough